OpenStudy (anonymous):

solve by elimination 0.3x-0.2y=4 0.4x+0.3y=-1 by the column method

OpenStudy (anonymous):

IS THIS ONE THE SIMULTANEOUS EQUATION ONE?

OpenStudy (anonymous):

i have no idea what that is

OpenStudy (anonymous):

im not sure but i worked it out and x=8.235...=140/17 y=-7.647..=-130/17 im not sure :S is that what ur after? probs not, but i gave it a go :)